OpenStudy (anonymous):

y is inversely proportional to the square of x and y=p for a particular value of x. Express the value of y in terms of p when this value of x is halved

jhonyy9 (jhonyy9):

y invers proportional to square of x you can write y = 1/x^2

jhonyy9 (jhonyy9):

than y=p and express the y value when the value of x is halved - so this mean y = 1/((x/2)^2) y = 1/((x^2)/4) y = 4/(x^2)

jhonyy9 (jhonyy9):

sandy - how you think that will be this in term of p presented ?

jhonyy9 (jhonyy9):

so we know that y = 1/(x^2) y=p so result p = 1/(x^2) so than y= 4/(x^2) in term of p will be y = 4p right ?

OpenStudy (anonymous):

Looks good.

jhonyy9 (jhonyy9):

thank you for your opinion good luck bye bye